# Iceberg Sensemaking: A Process Model for Critical Data Analysis

### v3

@@ 1 Introduction
Data analysis is the bedrock of many academic fields, but the responsible use of data means confronting the messy politics of knowledge. Critical studies of data, algorithms, and visualization have grown into a vital interdisciplinary field challenging the authority of data-driven systems and calling attention to the myriad ways they fail [1, 2, 3, 4, 5, 6, 7, 8]. Especially troubling cases of algorithmic bias have been tied to search engines [9], targeted advertising [10], predictive policing [11, 12], parole decisions [13], loan approval [14], and the general reinforcement of existing human prejudice under the guise of mechanical neutrality [15, 16]. AI-driven systems place data at an even further remove from human analysts, further reducing algorithmic transparency and galvanizing calls for a critical, humanistic understanding of data-driven reasoning and its limits.
@@
Tracing the source of these critiques leads deep into the philosophical underpinnings of the many fields and industries that increasingly rely on data to make sense of complicated subjects [17, 18]. With increasing reliance on data comes a tendency to adopt a basic, even tacit theory of knowledge rooted in positivism: the scientific stance that objective truth is accessible through experimental methods based in empirical measurement, and that proper application of these methods will yield results that are independent from the analyst’s own perspective and position in the world. Although the basic tenets of positivism are rarely stated aloud, a positivist stance toward knowledge and methodology predominates in visualization research and many of the fields where we collaborate with domain experts [19, 20]. A clear majority of visualization researchers are at least implicitly positivist in treating large-scale analysis as though it can minimize the subjective dimensions of human judgment. Sometimes this implicit positivism is practical and expedient as a default research mode, but it would be an error to imagine that this is how things must be done, especially in light of the mounting evidence that a positivist approach to data can often lead us astray.
@@
Objections to positivism are often grounded in interpretivism, the philosophical stance that knowledge is created through acts of interpretation, that many different interpretations of a given subject are possible, and the viability of different interpretations can be judged through critical examination of their assumptions, cogency, ethical implications, and explanatory power. For an interpretivist, you cannot separate data from human values, contexts, and practices because doing so is an erasure of essential complexities that shape data and knowledge alike. To resurface these erasures, interpretivist critiques often examine structural power, ideology, and other matters taken for granted as the natural order of things. When aimed at data analysis, the ultimate thrust of interpretivist critique is to see the separation of data from the politics of human knowledge as a fundamental misunderstanding of data itself.
@@
Our work in this paper is directed at the sensemaking process, a way of modeling the search for new understanding of a subject by gathering data and analyzing it to answer task-specific questions. Several influential sensemaking models have been proposed in fields such as HCI, statistics, and visual analytics [21, 22, 23], but these models evince the typical positivist tendency to minimize the challenging entanglement of human judgment in social, cultural, and political complexities. Moreover, these models focus on successful sensemaking and largely avoid questions about how sensemaking fails and why. Our core contribution is a new model for sensemaking that is built on interpretivist principles. The Iceberg Sensemaking model starts with the idea that data itself is not the fundamental basis of knowledge, but rather the tip of a larger iceberg constituting frameworks of knowledge, often called ‘schemas’ in the sensemaking literature. As a secondary contribution, we analyze the philosophical baggage inherent in the very concept of a schema, which has been largely unexamined in the sensemaking literature despite serving as a linchpin of past models.
@@
A second feature of our model is to place Schemas First and Always in sensemaking. Everyone carries tacit schemas into the sensemaking process through past experience, social conditioning, and other factors that guide our interpretation of a given problem. Schemas come into play before even a scrap of data is considered, and these schemas exist throughout the sensemaking process as vital components at every stage. This stance is a key point of distinction with current sensemaking models, where schemas are often treated like they enter the process only after some sensemaking work has already taken place.
The third feature of our model is to treat Data as a Schematic Artifact: every dataset rests on top of a schema that serves as the foundation for the seemingly raw facts and figures it contains. To deploy a familiar analogy, we treat data as the tip of the iceberg in sensemaking. A dataset's explicit schema is visible just under the waterline, while the tacit schema that guides its original creation and later interpretation is a submerged mass concealed in the deep waters of structural, hegemonic, disciplinary, and interpersonal power [27]—until and unless we go to the trouble of exploring it. Treating a dataset as objective truth rather than a designed artifact constitutes a failure to map these tacit schemas and thus ignore the schematic structure that has already given shape to it. When schemas are explicitly acknowledged and well-mapped, many pitfalls of data analysis can be confronted and addressed. Any time an analyst goes to the trouble of scrutinizing the assumptions underlying a particular dataset, they are already treating data as a schematic artifact.
@@ 2 Related Work
A process model classifies actions according to a set of stages, offering practical guidance on the order and transition criteria for progressing through those stages [28]. A sensemaking model is a type of process model that centers the human analyst in the act of gathering and interpreting information, capturing the “how” of a data analysis workflow, breaking it down into stages, and outlining the transitions that lead to a conclusion. We now discuss three major sensemaking models proposed in previous work. These summaries focus on the role of schemas and related concepts in each model.
@@ 2.2 Sensemaking in Data Science, Statistics, and Visualization
Another highly influential model of the sensemaking process covering data science is given by the statisticians Grolemund & Wickham [22]. Their model (Fig. 2) describes the sensemaking process through a parsimonious set of loops in which schemas, insights, and information (data) are the central elements. Grolemund & Wickham break down the process of data analysis into an exploratory stage and a confirmatory stage. Exploratory data analysis begins with information, then the analyst searches for a relevant schema to explain it. Confirmatory data analysis begins with a schema, followed by a search for data to confirm it. This distinction draws on the statistician John Tukey's [30, 31] seminal models of the data analysis process, which remain central to many modern frameworks for interactive data analysis [32]. Grolemund & Wickham define a schema as “a mental model that contains a breadth of information about a specific type of object or concept” (p.5). As we illustrate below, the key distinction between this model and ours is that Grolemund & Wickham treat schemas merely as mental models and not an active component of the data they work with, postponing the consideration of schemas to the point when the analyst constructs one explicitly.
@@ 2.3 Sensemaking in Visual Analytics
Sacha et al. [23] take a visual analytics approach to sensemaking (Fig. 3) that builds upon ideas from Pirolli & Card to place the interplay between visualization, models, and data as interrelated means of constructing knowledge for the human analyst. Their model also draws upon Norman's [33] gulfs of evaluation and execution, defining stages of interactions in which goals lead the analyst from hypotheses to actions (execution) and from findings to insights (evaluation). Their use of the word 'knowledge' corresponds to the role of 'schema' in the other two sensemaking models, and they define it simply as 'justified belief'. By treating knowledge (schema) as both the beginning and end point of the sensemaking process, this model is distinct from the previous two because it underlines that human understanding guides every stage of sensemaking. As we later discuss in greater detail, the major limitation of the Sacha model is its under-theorization of knowledge. Neither is ’justified belief’ an adequate definition, nor is the duality of tacit and explicit knowledge addressed.
@@ 3 Background: Understanding Schemas
Although the term schema is a standard feature of many sensemaking models, it is rarely defined, let alone given close attention as a word with a complicated history and several distinct meanings. We present a brief keyword study [34] examining how people have thought about schemas and related concepts over time. This broad examination will help delineate different stances toward schemas in philosophy, psychology, and critical theory, calling particular attention to the intersection of these theories with HCI and visualization design. We also briefly describe the use of the word ‘schema’ in the computer science database literature, mainly to distinguish the conventional sense of a database schema from the mental and social schemas that are more fundamental to the sensemaking process. This review spans a substantial historical and disciplinary range, answering a recent provocation from Meyer and Dykes [20] for visualization researchers to give greater attention to the philosophical groundwork underlying the field.
We group traditional theories of knowledge pertaining to schemas into two broad categories: cognitivist and interpretivist. The principal differences between these perspectives can be summarized as follows. For cognitivists, schemas are discrete, transportable structures of understanding that we acquire in the course of learning and finding order in the world. As we will show, the cognitivist approach to schemas in the sensemaking literature implicitly adopts a positivist theory of knowledge without acknowledging this paradigmatic baggage. The alternative we offer is interpretivism, which entails viewing schemas as something we construct and acquire through social practices, systems, and institutions that offer a shared means of framing things in the world, albeit one that is highly contingent and contestable. While it is beyond the scope of this article to survey every major touchstone in these disparate theories of knowledge, our focus is to delineate a divergence in the treatment of schemas and articulate an alternative path in visualization and adjacent fields.
In contrast to these two ways of thinking about schemas as mental structures, the word ‘schema’ has a much narrower technical definition in the computer science database literature. A database schema is a model describing relationships between entities within a particular dataset or database [35]. In data analysis, this kind of schema can be explicitly communicated through a data dictionary packaged with a dataset, or even through meaningful column headers in tabular spreadsheets. Although this definition is fairly distant from the philosophical concerns outlined below, it carries a historical and procedural connection to these other uses: a database schema is a technical instantiation of a mental schema, much like a schematic drawing in architecture or engineering.
@@ 3.1 Cognitive Schemas
The sense of the word schema as a mental construct was coined by the philosopher Immanuel Kant, who borrowed it from an ancient Greek word meaning shape or figure. In his Critique of Pure Reason [36], Kant described schemas as innate forms that shape the way perceptible phenomena appear to us. Schemas serve a crucial role in Kant’s elaborate theory of knowledge, mediating between our empirical sensation of the world and ideal forms that exist only in our minds. For instance, Kant treats time as a schema because it is neither an inherent property of the world nor an abstract concept like a triangle, but rather a mental construct that fundamentally orders our understanding of things. While this origin story may seem distant, it usefully illustrates how the concept of the schema was introduced into both philosophy and modern science: schemas are ordering principles that easily go unnoticed because they serve as the starting point for knowing anything at all.
The term schema was revived in the early twentieth century by the psychologist Jean Piaget, whose cognitivist theory of mind treats schemas as basic units of understanding that we acquire through inherent cognitive faculties[37]. For Piaget, schemas enter human cognition at the developmental stage when a child first grasps the permanence of objects. Piaget reasoned that something abstract (a mental schema) must be enabling the child to think about the object in its absence. This theory forms the initial basis of cognitive constructivism in psychology, holding that we acquire schemas over time through basic cognitive faculties and developmental processes. The fundamental power of this theory is its capacity to explain how a child apparently begins as a blank slate and develops into a wielder of various symbols, concepts, models, and categories. This cognitive approach to schemas influenced later models of the human mind derived from cybernetics [38, 39], logical positivism [40, 41, 42], formal linguistics [43], and analytic philosophy of mind [44, 45] to form the interdisciplinary field of cognitive science, where schemas are treated as stable, objectively comprehensible mechanisms of understanding. For a cognitivist, once someone learns the schema for a histogram, truth table, or road map, these are schemas one shares with everyone else who knows them. Nativists take an even stronger view, treating many schemas of perception, reasoning, and even moral judgment as matters of inborn knowledge [46].
@@
The application of cognitive science to human-computer interaction (HCI) and visualization led to many notable successes and technical breakthroughs even in these early years [53, 54, 55]. From 1970 forward, some of the most productive interdisciplinary exchanges between psychologists and computer scientists took place at Xerox PARC, where cognitivism informed fundamental R&D work on computer interfaces [56]. HCI researchers including Shneiderman [57] and Hutchins, Hollan, and Norman [58] recognized the benefits of direct manipulation of computers and often described their findings using a theoretical lens that included cognitive schemas. The foundational work in visualization from PARC closely adhered to a cognitivist framework, as laid out in the Cognitive Coprocessor Architecture [59]. This approach was followed by the vast majority of early information visualization researchers [60]. Another stream of psychology research involving schemas originates with Gibson [61], who departed from the strictly computational view of cognition with his theory of affordances, drawing attention toward the ways a subject may perceive or fail to perceive different uses of things in their environment. Norman (1988) made fundamental contributions to HCI by adapting Gibson's theory of affordances to design practices, describing the gulf of evaluation and gulf of execution as the key problems to be addressed by designers. Still, Gibson and Norman present a fairly positivist view of affordances as fundamental properties of objects rather than active constructions of their users.
@@
Although visualization as a field has matured and broadened over time, the cognitivist mentality has become the default for development and evaluation of new visual encoding and interaction techniques. With cognitivism has come the historical baggage of positivism, even though many visualization researchers do not identify as positivists. A recent survey of effective methods in visual data communication [62] offers a useful index of progress in this field while also displaying the tendency to view schemas through a purely cognitive lens. Upon presenting a graphic that’s just an ambiguous assortment of circles, the authors write: “If you are having trouble extracting the data from this visualization, it is not your fault—you do not have the needed schema” (2021, p. 132). This explanation succinctly illustrates the role schemas serve in many cognitive accounts of data visualization today: schemas are treated as discrete, transportable skills that you either have at your disposal, or don’t. Uses of a particular schema by different individuals are treated as equivalent, which is undeniably practical in many research scenarios, but comes at the expense of examining the complex social, cultural, and political factors underlying the acquisition, significance, limitations, and variations of these schemas when making sense of data.

### v4

@@ 1 Introduction
Data analysis is the bedrock of many academic fields, but the responsible use of data means confronting the messy politics of knowledge. Critical studies of data, algorithms, and visualization have grown into a vital interdisciplinary field challenging the authority of data-driven systems [1, 2, 3, 4, 5] and calling attention to the many ways these systems fail, skew their subjects, and obscure their inner workings [6, 7, 8]. Especially troubling cases of algorithmic bias have been tied to search engines [9], targeted advertising [10], predictive policing [11, 12], parole decisions [13], loan approval [14], and the general reinforcement of existing human prejudice under the guise of mechanical neutrality [15, 16]. AI-driven systems place data at an even further remove from human analysts, further reducing algorithmic transparency and motivating calls for critical, humanistic understanding of data-driven reasoning and its limits.
@@
Tracing the source of these critiques leads deep into the philosophical underpinnings of the many fields and industries that increasingly rely on data to make sense of complicated subjects [17, 18]. With increasing reliance on data comes a tacit tendency to adopt a theory of knowledge rooted in positivism: the scientific stance that objective truth is accessible through experimental methods based in empirical observation, and that proper application of these methods will yield results that are independent from the analyst’s own perspective and position in the world. Although the basic tenets of positivism are rarely stated aloud, a positivist stance toward knowledge and methodology predominates in visualization research and many of the fields where we most often collaborate with domain experts [19, 20]. A clear majority of visualization researchers are at least implicitly positivist in treating large-scale analysis as though it can minimize the subjective dimensions of human judgment. Sometimes this implicit positivism is practical and expedient as a default research mode, but it would be an error to imagine that this is how things must be done.
@@
Objections and alternatives to positivism are often grounded in interpretivism, the philosophical stance that knowledge is created through acts of interpretation, that many different interpretations of a given subject are viable, and the relative viability of different interpretations can be judged through critical examination of their assumptions, cogency, ethical implications, and explanatory power. For an interpretivist, you cannot separate data from human values, contexts, and practices because doing so is an erasure of essential complexities that shape data and knowledge alike. To resurface these erasures, interpretivist critiques often examine structural power, ideology, and other social phenomena that may otherwise be taken for granted as the natural order of things. To be clear, we are not talking about statistical power, but the power enmeshed in structures of society and politics. When aimed at data analysis, the ultimate thrust of interpretivist critique is to see the separation of data from the politics of human knowledge as a fundamental misunderstanding of data itself.
@@
Our work in this paper is directed at the sensemaking process, a way of modeling the search for new understanding of a subject by gathering data and analyzing it to answer task-specific questions. Although visualization is heavily used in sensemaking, matters of visual encoding and interaction design fall outside the scope of the sensemaking process. Several influential sensemaking models have been proposed in fields such as HCI [21], statistics [22], and visual analytics [23], but these models follow the typical positivist tendency to minimize how much of human judgment is shaped by social, cultural, and political factors. Moreover, these models focus on successful sensemaking and largely avoid questions about how sensemaking fails and why. Our core contribution is a new model for sensemaking that is built on interpretivist principles. The Iceberg Sensemaking model starts with the idea that data itself is not the fundamental basis of knowledge, but rather the tip of a larger iceberg constituting frameworks of knowledge, which are often called ‘schemas’ in the sensemaking literature. As a secondary contribution, we analyze the philosophical baggage inherent in the very concept of a schema, which has been largely unexamined in the sensemaking literature despite serving as a linchpin of past models.
@@
A second feature of our model is to place Schemas First and Always in sensemaking. Everyone carries tacit schemas into the sensemaking process through past experience, social conditioning, and other factors that guide interpretation of a given problem. Schemas come into play before even a scrap of data is considered, and these schemas exist throughout the sensemaking process as vital components at every stage. This stance is a key point of distinction with current sensemaking models, where schemas are often treated like they enter the process only after some sensemaking work has already taken place.
The third feature of our model is to treat Data as a Schematic Artifact: every dataset rests on top of a schema that serves as the foundation for the seemingly raw facts and figures it contains. To deploy a familiar analogy, we treat data as the tip of the iceberg in sensemaking. A dataset’s explicit schema is visible just above the waterline, while the tacit schema that guides its original creation and later interpretation is a submerged mass concealed in the deep waters of structural, hegemonic, disciplinary, and interpersonal power [27]—until and unless we go to the trouble of exploring it. Treating a dataset as objective truth rather than a designed artifact constitutes a failure to map these tacit schemas and thus ignore the schematic structure that has already given shape to the data. When schemas are explicitly acknowledged and well mapped, many pitfalls of data analysis can be confronted and addressed. Any time an analyst goes to the trouble of scrutinizing the assumptions underlying a particular dataset, they are already treating data as a schematic artifact.
@@ 2 Related Work
A process model classifies actions according to a set of stages, offering practical guidance on the order and transition criteria for progressing through those stages [28]. A sensemaking model is a type of process model that centers the human analyst in the act of gathering and interpreting information, capturing the “how” of a data analysis workflow, breaking it down into stages, and outlining the transitions that lead to a conclusion. We now summarize three major sensemaking models proposed in previous work. These summaries focus on the role of schemas and related concepts in each model.
@@ 2.2 Sensemaking in Data Science, Statistics, and Visualization
Another highly influential model of the sensemaking process covering data science is given by the statisticians Grolemund & Wickham [22]. Their model (Fig. 2) describes the sensemaking process through a parsimonious set of loops in which schemas, insights, and information (data) are the central elements. Grolemund & Wickham break down the process of data analysis into an exploratory stage and a confirmatory stage. Exploratory data analysis begins with information, then the analyst searches for a relevant schema to explain it. Confirmatory data analysis begins with a schema, followed by a search for data to confirm it. This distinction draws on the statistician Tukey’s [30, 31] seminal models of the data analysis process, which remain central to many modern frameworks for interactive data analysis [32]. Grolemund & Wickham define a schema as “a mental model that contains a breadth of information about a specific type of object or concept” (p.5). As we illustrate below, the key distinction between this model and ours is that Grolemund & Wickham treat schemas merely as mental models and not an active component of the data they work with, postponing the consideration of schemas to the point when the analyst constructs one explicitly.
@@ 2.3 Sensemaking in Visual Analytics
Sacha et al. [23] take a visual analytics approach to sensemaking (Fig. 3) that builds upon ideas from Pirolli & Card to place the interplay between visualization, models, and data as interrelated means of constructing knowledge for the human analyst. Their model also draws upon Norman’s [33] gulfs of evaluation and execution, defining stages of interactions in which goals lead the analyst from hypotheses to actions (execution) and from findings to insights (evaluation). Their use of the word ‘knowledge’ corresponds to the role of schemas in the other two sensemaking models, and they define it simply as ‘justified belief’. By treating knowledge (schema) as both the beginning and end point of the sensemaking process, this model is distinct from the previous two because it underlines that human understanding guides every stage of sensemaking. As we later discuss in greater detail, the major limitation of the Sacha model is its under-theorization of knowledge. ’Justified belief’ is not an adequate definition, nor is the duality of tacit and explicit knowledge addressed.
@@ 3 Background: Understanding Schemas
Although the term ‘schema’ is a standard feature of many sensemaking models, it is rarely defined, let alone given close attention as a word with a complicated history and several distinct meanings. We present a brief keyword study [34] examining how people have thought about schemas and related concepts over time. This broad examination will help delineate different stances toward schemas in philosophy, psychology, and critical theory, calling particular attention to the intersection of these theories with HCI and visualization design. We also briefly describe the use of the word ‘schema’ in the computer science database literature, mainly to distinguish the conventional sense of a database schema from the mental and social schemas that are more fundamental to the sensemaking process. This review spans a substantial historical and disciplinary range, answering a recent provocation from Meyer and Dykes [20] for visualization researchers to give greater attention to the philosophical groundwork underlying the field.
We group traditional theories of knowledge pertaining to schemas into two broad categories: cognitivist and interpretivist. The principal differences between these perspectives can be summarized as follows. For cognitivists, schemas are discrete, transportable structures of understanding that we acquire in the course of learning and finding order in the world. Cognitivism is the scientific domain from which visualization, as a field, has adopted many of its positivist tendencies. As we will show in the keyword study below, the cognitivist approach to schemas in the sensemaking literature implicitly adopts a positivist theory of knowledge without acknowledging this paradigmatic baggage. The alternative we offer is interpretivism, which entails viewing schemas as something we construct and acquire through social practices, systems, and institutions that offer a shared means of framing things in the world, albeit one that is highly contingent and contestable. While it is beyond the scope of this article to survey every major touchstone in these disparate theories of knowledge, our focus is to delineate a divergence in the treatment of schemas and articulate an alternative path in visualization and adjacent fields.
In contrast to these two ways of thinking about schemas as mental structures, the word ‘schema’ has a much narrower technical definition in the computer science database literature. A ‘database schema’ is a model describing relationships between entities within a particular dataset or database [35]. In data analysis, this kind of schema can be explicitly communicated through data dictionaries packaged with datasets, or even through meaningful column headers in tabular spreadsheets. Although this definition is far from the philosophical concerns outlined below, it carries a historical and procedural connection to these other uses: a database schema is a technical instantiation of a mental schema, much like a schematic drawing in architecture or engineering.
@@ 3.1 Cognitive Schemas
The sense of the word ‘schema’ as a mental construct was coined by the philosopher Immanuel Kant, who borrowed it from an ancient Greek word meaning shape or figure. In his Critique of Pure Reason [36], Kant described schemas as innate forms that determine the way perceptible phenomena appear to us. Schemas serve a crucial role in Kant’s elaborate theory of knowledge, mediating between our empirical sensation of the world and ideal forms that exist only in our minds. For instance, Kant treats time as a schema because it is neither an inherent property of the world nor an abstract concept like a triangle, but rather a mental construct that fundamentally orders our understanding of things. While this origin story may seem historically distant, it usefully illustrates how the concept of the schema was introduced into modern philosophy and science: schemas act as ordering principles that easily go unnoticed because they serve as the starting point for knowing anything at all.
The term ‘schema’ was revived in the early twentieth century by the psychologist Jean Piaget, whose cognitivist theory of mind treats schemas as basic units of understanding that we acquire through inherent cognitive faculties[37]. For Piaget, schemas enter human cognition at the developmental stage when a child first grasps the permanence of objects. Piaget reasoned that something abstract (a mental schema) must be enabling the child to think about the object in its absence. This theory forms the initial basis of cognitive constructivism in psychology, holding that we acquire schemas over time through basic cognitive faculties and developmental processes. The fundamental power of this theory is its capacity to explain how a child apparently begins as a blank slate and develops into a wielder of various symbols, concepts, models, and categories. This cognitive approach to schemas influenced later models of the human mind derived from cybernetics [38, 39], logical positivism [40, 41, 42], formal linguistics [43], and analytic philosophy of mind [44, 45] to form the interdisciplinary field of cognitive science, where schemas are treated as stable, objectively comprehensible mechanisms of understanding. For a cognitivist, once someone learns the schema for a histogram, truth table, or road map, these are schemas one shares with everyone else who knows them. Nativists take an even stronger view, treating many schemas of perception, reasoning, and even moral judgment as matters of inborn knowledge [46].
@@
The application of cognitive science to human-computer interaction (HCI) and visualization led to many notable successes and technical breakthroughs even in these early years [53, 54, 55]. From 1970 forward, some of the most productive interdisciplinary exchanges between psychologists and computer scientists took place at Xerox PARC, where cognitivism informed fundamental R&D work on computer interfaces [56]. HCI researchers including Shneiderman [57] and Hutchins, Hollan, and Norman [58] recognized the benefits of direct manipulation of computers and often described their findings using a theoretical lens that included cognitive schemas. The foundational work in visualization from PARC closely adhered to a cognitivist framework, as laid out in the Cognitive Coprocessor Architecture [59]. This approach was followed by the vast majority of early information visualization researchers [60]. Another stream of psychology research involving schemas originates with Gibson [61], who departed from the strictly computational view of cognition with his theory of affordances, drawing attention toward the ways a subject may perceive or else fail to perceive different uses of things in their environment. Norman [62, 33] made fundamental contributions to HCI by adapting Gibson’s theory of affordances to design practices, describing the gulf of evaluation and gulf of execution as the key problems to be addressed by designers. By viewing affordances as fundamental properties of objects rather than active constructions of their users, Gibson and Norman evince a tacit positivism drawn from the many cognitivist influences on early HCI and Design research.
@@
Although visualization as a field has matured and broadened over time, this cognitivist mentality has become the default for development and evaluation of new visual encoding and interaction techniques. Even though many visualization researchers do not identify as positivists, the field has inherited a strong inclination for positivist methods, values, and assumptions through the influence of cognitive psychology. A recent survey of effective methods in visual data communication [63] offers a useful index of progress in this field while also displaying the tendency to view schemas through a purely cognitive lens. Upon presenting a graphic that is just an ambiguous assortment of circles, the authors write: “If you are having trouble extracting the data from this visualization, it is not your fault—you do not have the needed schema” (2021, p. 132). This explanation succinctly illustrates the role schemas serve in many cognitive accounts of data visualization today: schemas are treated as discrete, transportable skills that you either have at your disposal, or don’t. The uses of a particular schema by different individuals are treated as equivalent, which is undeniably practical in many research scenarios, but comes at the expense of examining the complex social, cultural, and political factors underlying the acquisition, significance, limitations, and variations of these schemas when making sense of data.